### v1

# On world-sheet S-matrix of NSR string in static gauge

On world-sheet S-matrix
of NSR string in static gauge
Arkady A. Tseytlin [^1] Also at the Institute for Theoretical and Mathematical Physics (ITMP) of MSU and Lebedev Institute.      and     Zihan Wang
Abdus Salam Centre for Theoretical Physics
Imperial College London, SW7 2AZ, UK
zihan.wang18@imperial.ac.uk
As was shown in arXiv:1203.1054, expanding the Nambu action near the "long string" vacuum in the static gauge one finds that the one-loop $2\to 2$ scattering amplitude of the $D-2=24$ transverse 2d fluctuations $X^{i}$ is given by a pure phase expression consistent with underlying integrability. Similar computation in the Green-Schwarz superstring was carried out in arXiv:2404.09658. Here we consider the case of the NSR or spinning string starting with its manifestly 2d covariant action given by $D$ scalar multiplets coupled to 2d supergravity. Integrating out the zweibein and gravitino fields is not straightforward so that a Nambu-like version of the spinning string action involving only the scalar coordinates and their 2d fermionic partners was not given in the past. We show how the elimination of the auxiliary fields can be done in the static gauge after a particular choice of the superconformal gauges on the fermions while keeping the transverse fields $X^{i}$ and $\psi^{i}$ off-shell. The resulting one-loop S-matrix for $X^{i}$ turns out, as expected, to be the same as in the GS superstring case. We also discuss similar actions for the heterotic string and a $T\bar{T}$ deformation of a free 2d scalar multiplet.
1 Introduction
Starting with the Nambu action in flat $D$-dimensional space-time one may expand it near the "long string" vacuum $X^{a}=\xi^{a}$ ($a=0,1$). In the static gauge one gets an interacting action for the $D-2$ physical 2d massless scalars $X^{i}$
$S=-T\int d^{2}\xi\,\sqrt{-\det(\eta_{ab}+\partial_{a}X^{i}\partial_{b}X^{i})}=-T\int d^{2}\xi\,\big{[}1+{1\over 2}\partial_{a}X^{i}\partial^{a}X^{i}+{\cal O}((\partial X)^{4})\big{]}\ .$ (1.1)
The resulting 4-point scattering amplitude of $X^{i}$ is straightforward to compute at the tree and one-loop [1, 2] (and 2-loop [3]) orders and it is found to be consistent with integrability of the underlying theory for $D=26$.
Similar one-loop computation of the bosonic amplitude can be done in the $D=10$ Green-Schwarz superstring case [4] (cf. also [5, 6]). This is done by starting with the GS generalization of the Nambu action (found by solving for the independent 2d metric $g_{ab}$) and then fixing the static gauge on bosonic coordinates $X^{A}=(\xi^{a},X^{i}(\xi))$ and an appropriate $\kappa$-symmetry gauge on their fermionic counterparts $\theta$. The resulting action expanded in terms of 8 commuting and 16 real anticommuting physical components of the fields can be used to compute the 4-scalar scattering amplitude containing the fermionic loop contribution.
One may wonder how an analogous computation of S-matrix in the long string vacuum can be done in the case of the NSR or spinning string that has manifest world-sheet supersymmetry. The corresponding bosonic amplitude is expected to be the same as the GS string one in $D=10$ (two theories are equivalent in the light-cone gauge [7, 8]). However, it turns out to be non-trivial to demonstrate this by starting directly with the covariant spinning string action found in [9, 10]. The corresponding Lagrangian describes $D$ copies of 2d real scalar multiplet $(X^{A},\psi^{A})$ coupled to 2d $N=1$ conformal supergravity fields $(e^{a}_{\mu},\chi_{\mu})$
$\mathcal{L}=$ $e\Big{[}-\frac{1}{2}\partial_{\mu}X^{A}\partial^{\mu}X_{A}-\frac{i}{2}\bar{\psi}^{A}\gamma^{\mu}\partial_{\mu}\psi_{A}+\frac{i}{2}\bar{\chi}_{\mu}\gamma^{\nu}\gamma^{\mu}\psi^{A}\partial_{\nu}X_{A}-\frac{1}{16}\bar{\chi}_{\mu}\gamma^{\nu}\gamma^{\mu}\chi_{\nu}\bar{\psi}^{A}\psi_{A}\Big{]},$ (1.2)
where $\gamma^{\mu}=e_{a}^{\mu}\gamma^{a}\,,\ \{\gamma^{a},\gamma^{b}\}=2\eta^{ab}$. A reparametrization-covariant Nambu-like counterpart of this action involving only $X^{A}$ and $\psi^{A}$ was not discussed in the past: given a non-linear dependence of (1.2) on the gravitino $\chi_{\mu}$ is not immediately clear how to integrate it out and then solve for $e^{a}_{\mu}$.
Fixing the static gauge $X^{a}=\xi^{a}$, to compute the scattering amplitude for $X^{i}$ one needs an action containing only the physical components $(X^{i},\psi^{i})$ that should not be constrained by any additional conditions. A seemingly natural approach is to use the local superconformal symmetry to gauge-fix $\chi_{\mu}=0$ and then try to solve the constraint resulting from variation over $\chi_{\mu}$ to eliminate the longitudinal components $\psi^{a}=(\psi^{0},\psi^{1})$ but this appears to be hard to implement explicitly.
Our aim below will be to show how this problem can be overcome using a different gauge fixing procedure. [^2] The computation of one-loop scattering amplitude for $X^{i}$ in the NSR case was previously discussed in [6] using an action for $(X^{i},\psi^{i})$ found by requiring a non-linear realization of the underlying $D$-dimensional Poincare symmetry. This approach did not start with a fully covariant spinning string action and it appears to be hard to verify its validity (e.g. the uniqueness of the resulting action). This will be the topic of section 2. In section 3 we will discuss similar static-gauge actions for the heterotic string and GS superstring. In section 4 we will compute the spinning string one-loop scattering amplitude for 4 massless scalars $X^{i}$ and show that the result is the same as in the superstring case in [4].
One motivation to study the action of the spinning string in the static gauge is to compare it to a $T\bar{T}$ deformation of the free 2d scalar multiplet $(X^{i},\psi^{i})$ (cf. [11, 12, 13, 14]). We will comment on this in section 2.4 below.
Both the Nambu and the GS string world-sheet S-matrix computations can be generalized to the bosonic membrane and space-time supersymmetric M2-brane [15] cases [16, 4]. One may wonder if a similar computation can be done for a 3d world-volume supersymmetric analog of the bosonic membrane action. Like in the spinning string case, to have 3d reparametrization invariance and supersymmetry one is to couple a collection of $D$ 3d real scalar multiplets $(X^{A},\psi^{A})$ to a 3d supergravity multiplet. Attempts to construct such 3d supersymmetric membrane action were discussed in [17, 18, 19]. In contrast to the 2d spinning string case where the Einstein term in the supergravity action is a total derivative this is not so in the 3d case and a consistent 3d supersymmetric analog of the membrane action appears to require the inclusion of the 3d supergravity part (cf. [20, 21]). In this case the metric and gravitino become dynamical and their contributions should be properly accounted for when computing the S-matrix for the bosonic $X^{i}$ fields. We hope to discuss this computation in the future.
2 Spinning string action expanded near long string vacuum
To compute the world-sheet S-matrix in the long string vacuum the first step is to expand the action (1.2) in the static gauge. This requires gauge fixing all local symmetries and integrating out the auxiliary fields to end up with an action for the physical $D-2$ components of the bosonic and fermionic fields $(X^{i},\psi^{i})$.
2.1 Gauge fixing
Let us first recall the notation we used in the spinning string action (1.2) (see also Appendix A). The world-sheet coordinates are $\xi^{a}$ ($a=0,1$), the induced metric is $g_{\mu\nu}=e_{\mu}^{a}e_{\nu}^{b}\eta_{ab},\ \eta_{ab}={\rm diag}(-1,1)$,   ($\mu,\nu=0,1$). The index $A=0,...,D-1$ labels the flat target space coordinates. $\psi^{A}$ are 2d Majorana fermions in real representation with $\bar{\psi}^{A}=(\psi^{A})^{T}\gamma^{0}$, and constant flat-space Dirac matrices ${\gamma}^{a}$ are related to ${\gamma}^{\mu}$ in (1.2) by $e^{a}_{\mu}$. We will denote the spinor components of $\psi$ as $\begin{pmatrix}\psi_{R}\\ \psi_{L}\end{pmatrix}$, with $\bar{\psi}=(\psi_{L},-\psi_{R})$ and similarly for $\chi_{\mu}$.
In addition to the reparametrization invariance and local Lorentz rotations given by
$\delta\psi^{A}=\frac{1}{2}\alpha\gamma^{*}\psi^{A},\qquad\delta e^{a}_{\mu}=\alpha\epsilon^{a}{}_{b}e^{b}_{\mu},\qquad\delta\chi_{\mu}=\frac{1}{2}\alpha\gamma^{*}\chi_{\mu}\,,\qquad\gamma^{*}\equiv{1\over 2}\epsilon_{ab}{\gamma}^{a}{\gamma}^{b}={\gamma}^{0}{\gamma}^{1}\ ,$ (2.1)
the action (1.2) is invariant under the local supersymmetry transformations (see [9, 10] for details)
$\delta X^{A}=i\bar{\varepsilon}\psi^{A}\,,\qquad\delta\psi^{A}=\gamma^{\mu}(\partial_{\mu}X^{A}-\frac{i}{2}\bar{\chi}_{\mu}\psi^{A})\varepsilon\,,\qquad\delta e^{a}_{\mu}={i}\bar{\varepsilon}\gamma^{a}\chi_{\mu}\,,\qquad\delta\chi_{\mu}=2\nabla_{\mu}\varepsilon\,,$ (2.2)
and also the Weyl and super-Weyl transformations
$\delta X^{A}=0\,,\quad\delta e^{a}_{\mu}=\lambda e^{a}_{\mu}\,,\quad\delta\psi^{A}=-\frac{1}{2}\lambda\psi^{A}\,,\quad\delta\chi_{\mu}=\frac{1}{2}\lambda\chi_{\mu},\qquad\qquad\delta\chi_{\mu}=\gamma_{\mu}\eta\,.$ (2.3)
We shall fix the reparametrizations by choosing the static gauge, i.e. setting the fluctuations of $X^{a}$ to zero,
$X^{A}=(\xi^{a},X^{i}),\qquad a=0,1,\ \ \ \qquad i=2,3,...,D-1.$ (2.4)
Then the variation under the local supersymmetry (2.2) of the first two components of $\psi^{A}$ takes the form $\delta\psi^{a}=\gamma^{\mu}\delta_{\mu}^{a}\varepsilon-\frac{i}{2}\bar{\chi}_{\mu}\psi^{a}\varepsilon$. The presence of the first inhomogeneous term implies that half of the $\psi^{a}$ components are Stueckelberg fields so that they can be eliminated by a gauge fixing. [^3] This parallels the logic that underlines the bosonic static gauge fixing: under the reparametrizations $\delta X^{a}=\zeta^{\mu}\partial_{\mu}X^{a}$ so that for $X^{a}=\xi^{a}+\tilde{X}^{a}$ one gets $\delta\tilde{X}^{a}=\zeta^{\mu}\delta^{a}_{\mu}+\zeta^{\mu}\partial_{\mu}\tilde{X}^{a}$. This implies that $\tilde{X}^{a}$ is a Stueckelberg field that can be gauge-fixed to zero. Explicitly, we shall fix the local supersymmetry by the following gauge [^4] From (2.2) one has $\delta(\gamma^{a}\psi_{a})=M\varepsilon$ where $M\equiv\gamma^{a}\gamma^{\rho}\big{(}\eta_{\rho a}-\frac{i}{2}\bar{\chi}_{\rho}\psi_{a}\big{)}$ is a diagonal 2 by 2 matrix in spinor components which is thus invertible, allowing one to find an $\varepsilon$ such that (2.5) is satisfied.
$\gamma_{a}\psi^{a}=0,\qquad\ \ \ \ i.e.\ \ \ \ \qquad\psi_{R}^{-}=0,\ \ \ \psi_{L}^{+}=0,\ \ \ \ \ \ \qquad\psi^{\pm}\equiv\psi^{0}\pm\psi^{1}\ .$ (2.5)
An important consequence of (2.5) is that the kinetic term for $\psi^{a}$ in the action (1.2) then vanishes (see (A.5) and (A.6))
$\bar{\psi}^{a}\gamma^{\mu}\partial_{\mu}\psi_{a}$ $=e^{\mu}_{b}\bar{\psi}^{a}\gamma^{b}\partial_{\mu}\psi_{a}=0.$ (2.6)
The super-Weyl transformation of the gravitino in (2.3) implies that $\gamma^{\mu}\chi_{\mu}$ is another Stueckelberg field that can be fixed by the gauge
$\gamma^{\mu}\chi_{\mu}=0\ .$ (2.7)
As a result, the Lagrangian in (1.2) takes the form
$\mathcal{L}=e\Big{[}-\frac{1}{2}$ $(\eta_{\mu\nu}+\partial_{\mu}X^{i}\partial_{\nu}X^{i})g^{\mu\nu}-\frac{i}{2}\bar{\psi}^{i}\gamma^{\mu}\partial_{\mu}\psi^{i}+{i}\bar{\chi}^{\mu}\psi_{\underline{\mu}}+{i}\bar{\chi}^{\mu}\psi^{i}\partial_{\mu}X^{i}-\frac{1}{8}\bar{\chi}^{\mu}\chi_{\mu}\bar{\psi}^{A}\psi_{A}\Big{]},$ (2.8)
where we introduced the notation
$\eta_{\mu\nu}\equiv\delta^{a}_{\mu}\delta^{b}_{\nu}\eta_{ab}\ ,\qquad\qquad\psi_{\underline{\mu}}\equiv\delta^{a}_{\mu}\psi_{a}\ .$ (2.9)
After imposing the gauge conditions (2.5) and (2.7) we are left with 2+2 independent spinor components of the fermionic fields $\psi^{a}$ and $\chi_{\mu}$. These enter (2.8) without derivatives and thus play the role of auxiliary fields that can be integrated out algebraically.
Similar decomposition into Stueckelberg parts (that can be gauge-fixed) and auxiliary parts (that should be integrated out) applies to the zweibein. As follows from (2.1) and (2.3) the Weyl factor and antisymmetric part of $e_{a\mu}\equiv\eta_{ab}e^{b}_{\mu}$ can be gauge-fixed, e.g., by setting $e_{a\mu}$ to be symmetric matrix with determinant one, i.e. [^5] As was noted in [10], by fixing the gauges $\gamma^{\mu}\chi_{\mu}=0$, $e^{0}_{1}=0$ and $e=1$ the action (1.2) becomes equivalent to the one constructed in [22, 23].
$e_{[a\mu]}=0,\ \ \qquad\ \ \ \quad e\equiv\det e^{a}_{\mu}=1\ ,$ (2.10)
while the remaining 2 independent components of $e^{a}_{\mu}$ should be integrated out.666 In general, the strategy should be first (i) gauge-fixing Stueckelberg fields and then (ii) integrating out (or solving for) the remaining auxiliary fields. As Stueckelberg fields transform via algebraic shifts by gauge parameters, there are no non-trivial ghost determinants associated with setting them to zero.
2.2 Solving for the auxiliary fields
Let us first consider eliminating the remaining components of $\psi^{a}$ and $\chi_{\mu}$. Since these auxiliary fields appear in the Lagrangian (2.8) without derivatives, the result of integrating them out is an "ultralocal" contribution $\sim\delta^{(2)}(0)$ that can be absorbed into the path integral measure or regularized away (it vanishes in dimensional regularization that we will use below).
Indeed, viewing the 4-fermion term $\bar{\chi}^{\mu}\chi_{\mu}\bar{\psi}^{A}\psi_{A}$ as an interaction vertex, one may use perturbation theory by expanding in powers of this vertex near the quadratic fermionic part given by $\chi\psi$ terms in (2.8). The contribution of the quadratic fermionic terms produces a local determinant while the "Feynman diagrams" generated by the quartic fermionic vertex will also be proportional to $\delta^{(2)}(0)$. As a result, we are left just with the $\chi$-independent part of (2.8), i.e. ($g^{\mu\nu}=e^{\mu}_{a}e^{\nu}_{b}\eta^{ab}$)
$\mathcal{L}=-\frac{1}{2}e\Big{[}$ $(\eta_{\mu\nu}+\partial_{\mu}X^{i}\partial_{\nu}X^{i})g^{\mu\nu}+ie^{\mu}_{a}\bar{\psi}^{i}\gamma^{a}\partial_{\mu}\psi^{i}\Big{]}\ .$ (2.11)
The same conclusion is reached also by just solving the non-linear classical equations for $\chi_{\mu}$ and $\psi^{a}$ that follow from (2.8). Taking variations of (2.8) over $\chi_{\mu}$ and $\psi_{a}$ subject to the gauge conditions (2.7) and (2.5) we get (we use the same notation as in (2.9))
$i\gamma^{\nu}\gamma^{\mu}\psi_{\underline{\nu}}$ $+i\gamma^{\nu}\gamma^{\mu}\psi^{i}\partial_{\nu}X^{i}-\frac{1}{2}\chi^{\mu}\bar{\psi}^{A}\psi_{A}=0,$ (2.12)
$i\gamma_{\underline{\rho}}\gamma^{a}\chi^{\rho}-\frac{1}{2}(\bar{\chi}_{\mu}\chi^{\mu})\psi^{a}=0\ .$ (2.13)
Eq. (2.13) implies that (cf. (2.5))
$\chi^{+}_{R}=-\frac{i}{4}(\bar{\chi}^{\mu}\chi_{\mu})\psi^{+}_{R}\ ,\quad\qquad\qquad\chi^{-}_{L}=-\frac{i}{4}(\bar{\chi}^{\mu}\chi_{\mu})\psi^{-}_{L},$ (2.14)
where $\chi^{\pm}=\chi^{0}\pm\chi^{1}$ with $\chi^{\mu}=(\chi^{0},\chi^{1}$) and $L$ and $R$ label the upper and lower spinor components. The gauge condition (2.7) implies that (here $e^{0}_{0}$, etc., are the components of $e^{a}_{\mu}$)
$\chi^{-}_{R}=-\frac{e^{0}_{0}+e^{0}_{1}-e^{1}_{0}-e^{1}_{1}}{e^{0}_{0}-e^{0}_{1}-e^{1}_{0}+e^{1}_{1}}\chi^{+}_{R}\ ,\quad\qquad\qquad\chi^{+}_{L}=-\frac{e^{0}_{0}-e^{0}_{1}+e^{1}_{0}-e^{1}_{1}}{e^{0}_{0}+e^{0}_{1}+e^{1}_{0}+e^{1}_{1}}\chi^{-}_{L}\ ,$ (2.15)
$\bar{\chi}^{\mu}\chi_{\mu}=\,F(e)\,\chi^{-}_{L}\chi^{+}_{R}\ ,\qquad\qquad\qquad\ \ F(e)\equiv-\frac{4e^{2}}{[(e^{0}_{0}+e^{1}_{1})^{2}-(e^{0}_{1}+e^{1}_{0})^{2}]}\ .$ (2.16)
Then (2.14) may be written as
$\chi^{+}_{R}\Big{(}1-\frac{i}{4}F(e)\,\chi^{-}_{L}\psi^{+}_{R}\Big{)}=0,\qquad\qquad\chi^{-}_{L}\Big{(}1+\frac{i}{4}F(e)\,\chi^{+}_{R}\psi^{-}_{L}\Big{)}=0\ ,$ (2.17)
which is solved by [^7] The expressions in the brackets in (2.17) are invertible and thus setting them to zero does not give extra solutions.
$\chi^{+}_{R}=\chi^{-}_{L}=0\ .$ (2.18)
Combined with (2.15) this gives $\chi_{\mu}=0$ leading to (2.11). Note that (2.12) with $\chi_{\mu}$ determines $\psi_{a}$ in terms of $\psi^{i}\partial X^{i}$ but this has no consequence as $\psi_{a}$ drops out of the action.
The next step is to eliminate from (2.11) the remaining two auxiliary fields represented by $e^{a}_{\mu}$ subject to the gauge conditions (2.10). This can be done by taking variation of (2.11) over $e^{a}_{\mu}$ constrained by (2.10) (i.e. adding the corresponding projectors), solving for $e^{a}_{\mu}$ in terms of $(X^{i},\psi^{i})$ and plugging the result back into the Lagrangian (2.11). Equivalently, we may add the gauge conditions (2.10) to (2.11) with Lagrange multipliers and vary unconstrained $e^{a}_{\mu}$.
The resulting variation over the zweibein leads to the condition of the vanishing of the effective traceless and symmetric part of the stress tensor corresponding to (2.11) [^8] An alternative to imposing the Lorentz and Weyl gauges (2.10) on $e^{a}_{\mu}$ in order to eliminate it from (2.11) is to redefine $\psi^{i}$ so that to make the Lagrangian (2.11) manifestly Lorentz and Weyl invariant. Indeed, rescaling $\psi^{i}$ by $e^{-1/4}$ we find the fermionic term in (2.11) taking the form $ie\,e^{\mu}_{a}\,\bar{\psi}^{i}\gamma^{a}\partial_{\mu}\psi^{i}\to ie^{1/2}e^{\mu}_{a}\psi^{\prime i}\gamma^{a}\partial_{\mu}\psi^{\prime i}$. The resulting analog (2.11) is now invariant under the rescaling of $e^{a}_{\mu}$ with $\psi^{\prime i}$ not transforming. The variation of the action over $e^{a}_{\mu}$ is then manifestly traceless (with no need to use the fermionic equations of motion). Similarly, we may Lorentz-rotate $\psi^{i}$ and $e_{a\mu}$ to make the latter symmetric so that the variation of the resulting (2.11) over the zweibein will also be manifestly symmetric. We then end up with the same relation as in (2.19).
$e^{\nu}_{(a}\big{(}\eta_{\mu)\nu}+\partial_{\mu)}X^{i}\partial_{\nu}X^{i}\big{)}+\frac{i}{2}\bar{\psi}^{i}\gamma_{(a}\partial_{\mu)}\psi^{i}-\frac{1}{2}e_{a\mu}\Big{[}\big{(}\eta_{\lambda\nu}+\partial_{\lambda}X^{i}\partial_{\nu}X^{i}\big{)}g^{\lambda\nu}+\frac{i}{2}e^{\lambda}_{b}\,\bar{\psi}^{i}\gamma^{b}\partial_{\lambda}\psi^{i}\Big{]}=0\ .$ (2.19)
Contracting (2.19) with $e^{a}_{\nu}$, symmetrizing the indices $\mu$ and $\nu$, and taking the determinant allows to rewrite the Lagrangian in (2.11) in the "square-root" form
$\mathcal{L}$ $=-\sqrt{-\det\Big{[}\eta_{\mu\nu}+\frac{i}{2}e^{a}_{\mu}e^{\rho}_{(a}\partial_{\nu)}X^{i}\partial_{\rho}X^{i}+\frac{i}{2}e^{a}_{\nu}e^{\rho}_{(a}\partial_{\mu)}X^{i}\partial_{\rho}X^{i}+\frac{i}{4}e^{a}_{\mu}\bar{\psi}^{i}\gamma_{(a}\partial_{\nu)}\psi^{i}+\frac{i}{4}e^{a}_{\nu}\bar{\psi}^{i}\gamma_{(a}\partial_{\mu)}\psi^{i}\Big{]}}$
$\qquad\qquad-\frac{i}{4}e^{\mu}_{a}\,\bar{\psi}^{i}\gamma^{a}\partial_{\mu}\psi^{i}\ .$ (2.20)
Here $e^{a}_{\mu}$ should be replaced by its expression in terms of $X^{i}$ and $\psi^{i}$ found by solving (2.19).
Eq. (2.19) is a non-linear algebraic equation on the matrix $e^{a}_{\mu}$ subject to (2.10). It does not seem possible to write its solution in a closed form but it is sufficient for our goals here to solve it perturbatively by expanding in powers of $X^{i}$ and $\psi^{i}$, i.e.
$e^{a}_{\mu}=\delta^{a}_{\mu}+{\tilde{e}}^{a}_{\mu}+...,\qquad\qquad g_{\mu\nu}=\eta_{\mu\nu}+2\eta_{a(\mu}{\tilde{e}}^{a}_{\nu)}+...\ ,$ (2.21)
$\tilde{e}_{a\mu}=\frac{1}{2}\partial_{a}X^{i}\partial_{\mu}X^{i}+\frac{i}{4}\bar{\psi}^{i}\gamma_{(a}\partial_{\mu)}\psi^{i}-{1\over 4}\eta_{a\mu}\eta^{\sigma\rho}\Big{[}\partial_{\sigma}X^{i}\partial_{\rho}X^{i}+\frac{i}{2}\bar{\psi}^{i}\gamma_{a}\delta^{a}_{(\sigma}\partial_{\rho)}\psi^{i}\Big{]}\ ,$ (2.22)
where we imposed the conditions (2.10). This procedure can be extended to higher orders of expansion of $e^{a}_{\mu}$ and the result should be plugged into (2.20).
The final step is to expand (2.20) in powers of derivatives of the fluctuations $(X^{i},\psi^{i})$ near the long string vacuum. To compute the tree-level and one-loop $2\rightarrow 2$ scattering amplitude of $X^{i}$ one needs to keep only the following quadratic and quartic terms [^9] Note that there are no cubic $\psi\psi X$ vertices. To compute the one-loop tadpole contribution to the amplitude requires also the knowledge of 6-point vertices but like in [16, 4] the resulting contribution vanishes in dimensional regularization.
$\mathcal{L}=$ $-\frac{1}{2}\partial_{a}X^{i}\partial^{a}X^{i}-{i}\bar{\psi}^{i}\gamma^{a}\partial_{a}\psi^{i}$ (2.23)
$+\frac{1}{4}\partial_{a}X^{i}\partial_{b}X^{i}\partial^{a}X^{j}\partial^{b}X^{j}-\frac{1}{8}\partial_{a}X^{i}\partial^{a}X^{i}\partial_{b}X^{j}\partial^{b}X^{j}$
$+\frac{i}{2}\partial_{a}X^{i}\partial_{b}X^{i}\bar{\psi}^{j}\gamma^{a}\partial^{b}\psi^{j}-\frac{i}{4}\partial_{a}X^{i}\partial^{a}X^{i}\bar{\psi}^{j}\gamma^{b}\partial_{b}\psi^{j}$
$-\frac{1}{4}\bar{\psi}^{i}\gamma_{a}\partial_{b}\psi^{i}\bar{\psi}^{j}\gamma^{(a}\partial^{b)}\psi^{j}+\frac{1}{8}\bar{\psi}^{i}\gamma^{a}\partial_{{a}}\psi^{i}\bar{\psi}^{j}\gamma^{b}\partial_{b}\psi^{j}+....$
Here we dropped the -1 constant, rescaled $\psi\rightarrow\sqrt{2}\psi$ and used $a,b$ for the flat-space indices.
3 Other similar actions
3.1 Heterotic string
The covariant heterotic string action [24] can be found by coupling $(1,0)$ supersymmetric 2d scalar multiplet $(X^{A},\psi^{A})$ (with $\psi^{A}$ being left MW spinors) to the corresponding supergravity multiplet $(e^{a}_{\mu},\chi_{\mu}$) (with $\chi_{\mu}$ given by right MW spinors) and adding also $N=32$ right MW spinors $\varphi^{\rm r}$ (${\rm r}=1,2,...,N$) coupled to $e^{a}_{\mu}$. The resulting Lagrangian can be found from (1.2) by adding the chiral projectors on $\psi^{A}$ and $\chi_{\mu}$ and adding the kinetic term for $\varphi^{\rm r}$, i.e.
$\mathcal{L}=$ $e\Big{[}-\frac{1}{2}\partial_{\mu}X^{A}\partial^{\mu}X_{A}-\frac{i}{2}\bar{\psi}^{A}\gamma^{\mu}\partial_{\mu}\psi_{A}+\frac{i}{2}\bar{\chi}_{\mu}\gamma^{\nu}\gamma^{\mu}\psi^{A}\partial_{\nu}X_{A}-\frac{i}{2}\bar{\varphi}^{{\rm r}}\gamma^{\mu}\partial_{\mu}\varphi^{{\rm r}}\Big{]},$ (3.1)
$P_{+}\psi^{A}=\psi^{A},\ \qquad P_{-}\chi_{\mu}=\chi_{\mu},\qquad P_{-}\varphi^{\rm r}=\varphi^{\rm r},\qquad P_{\pm}\equiv\frac{1}{2}(1\pm\gamma^{*})\ .$ (3.2)
In the ${\gamma}^{a}$-matrix representation we are using the chiral fermions will have one of the two spinor components set to zero, $\psi^{A}_{R}=0$, $\chi_{L\mu}=0$, $\varphi^{\rm r}_{L}=0$. The corresponding action is invariant under local $(1,0)$ supersymmetry, i.e. (2.2) with the parameter subject to $P_{-}\varepsilon=\varepsilon$, i.e. $\varepsilon_{L}=0$ (see, e.g., [25]).
To expand (3.1) in the static gauge we may fix the supersymmetry and superconformal gauges as in (2.5),(2.7). The absence of the quadratic gravitino term in (3.1) simplifies its elimination from the action and we end up with the following analog of (2.11)
$\mathcal{L}=-\frac{1}{2}e\Big{[}(\eta_{\mu\nu}+\partial_{\mu}X^{i}\partial_{\nu}X^{i})g^{\mu\nu}+ie^{\mu}_{a}\bar{\Psi}^{{\mathcal{I}}}\gamma^{a}\partial_{\mu}\Psi^{{\mathcal{I}}}\ \Big{]}\ ,$ (3.3)
$\bar{\Psi}^{{\mathcal{I}}}\gamma^{a}\partial_{\mu}\Psi^{{\mathcal{I}}}\equiv\bar{\psi}^{i}P_{-}\gamma^{a}\partial_{\mu}\psi^{i}+\bar{\varphi}^{\rm r}P_{+}\gamma^{a}\partial_{\mu}\varphi^{\rm r}\ .$ (3.4)
To integrate out $e^{a}_{\mu}$ we may fix again the Lorentz and Weyl gauges as in (2.10) ending with a direct analogs of the equation (2.19) and the Lagrangian (2.20). Expanding it in powers of derivatives of all fields we get the following counterpart of (2.23)
$\mathcal{L}=$ $-\frac{1}{2}\partial_{a}X^{i}\partial^{a}X^{i}-{i}\bar{\Psi}^{{\mathcal{I}}}\gamma^{a}\partial_{a}\Psi^{{\mathcal{I}}}$ (3.5)
$+\frac{1}{4}\partial_{a}X^{i}\partial_{b}X^{i}\partial^{a}X^{j}\partial^{b}X^{j}-\frac{1}{8}\partial_{a}X^{i}\partial^{a}X^{i}\partial_{b}X^{j}\partial^{b}X^{j}$
$+\frac{i}{2}\partial_{a}X^{i}\partial_{b}X^{i}\bar{\Psi}^{{\mathcal{I}}}\gamma^{a}\partial^{b}\Psi^{{\mathcal{I}}}-\frac{i}{4}\partial_{a}X^{i}\partial^{a}X^{i}\bar{\Psi}^{{\mathcal{I}}}\gamma^{b}\partial_{b}\Psi^{{\mathcal{I}}}$
$-\frac{1}{4}\bar{\Psi}^{{\mathcal{I}}}\gamma_{a}\partial_{b}\Psi^{{\mathcal{I}}}\bar{\Psi}^{{\mathcal{J}}}\gamma^{(a}\partial^{b)}\Psi^{{\mathcal{J}}}+\frac{1}{8}\bar{\Psi}^{{\mathcal{I}}}\gamma^{a}\partial_{{a}}\Psi^{{\mathcal{I}}}\bar{\Psi}^{{\mathcal{J}}}\gamma^{b}\partial_{b}\Psi^{{\mathcal{J}}}+...\ .$
3.2 GS superstring
It is instructive also to compare (2.23) to a similar expansion of the GS action near the long string vacuum [4]. Solving for the independent metric $g_{\mu\nu}$ the type IIA GS superstring Lagrangian may be written as the sum of the "volume" and WZ terms [26]
$L=L_{1}+L_{2}\ ,\qquad\qquad L_{1}=-\sqrt{-g},$ (3.6)
$L_{2}=-{i}\epsilon^{\mu\nu}\partial_{\mu}X^{A}\Big{[}\bar{\theta}^{1}\Gamma_{A}\partial_{\nu}\theta^{1}-\bar{\theta}^{2}\Gamma_{A}\partial_{\nu}\theta^{2}\Big{]}+\epsilon^{\mu\nu}\bar{\theta}^{1}\Gamma^{A}\partial_{\mu}\theta^{1}\bar{\theta}^{2}\Gamma_{A}\partial_{\nu}\theta^{2},$ (3.7)
$g_{\mu\nu}=\Pi_{\mu}^{A}\Pi_{A\nu}\ ,\qquad\qquad\Pi_{\mu}^{A}=\partial_{\mu}X^{A}-i\bar{\theta}^{I}\Gamma^{A}\partial_{\mu}\theta^{I},$ (3.8)
where $\theta^{I}$ ($I=1,2$) are the 10d MW spinors of opposite chirality. Fixing the static gauge $X^{a}=\xi^{a}$ and expanding in derivatives we get
$L_{1}=$ $-\sqrt{-\det\Big{[}\eta_{\mu\nu}+\partial_{\mu}X^{i}\partial_{\nu}X^{i}-2i\bar{\theta}^{I}\Gamma_{(\mu}\partial_{\nu)}\theta^{I}+...\Big{]}}$
$=$ $-1-\frac{1}{2}\partial_{\mu}X^{i}\partial^{\underline{\mu}}X^{i}+i\bar{\theta}^{I}\Gamma_{\mu}\partial^{\underline{\mu}}\theta^{I}$
$+\frac{1}{4}\partial_{\mu}X^{i}\partial_{\nu}X^{i}\partial^{\underline{\nu}}X^{j}\partial^{\underline{\mu}}X^{j}-i\bar{\theta}^{I}\Gamma_{(\underline{\mu}}\partial_{\nu)}\theta^{I}\partial^{\underline{\nu}}X^{j}\partial^{\underline{\mu}}X^{j}-\bar{\theta}^{I}\Gamma_{(\underline{\mu}}\partial_{\nu)}\theta^{I}\bar{\theta}^{J}\Gamma^{\underline{\mu}}\partial^{\underline{\nu}}\theta^{J}$
$-\frac{1}{8}\partial_{\mu}X^{i}\partial^{\underline{\mu}}X^{i}\partial_{\nu}X^{j}\partial^{\underline{\nu}}X^{j}+\frac{i}{2}\bar{\theta}^{I}\Gamma^{\underline{\mu}}\partial_{{\mu}}\theta^{I}\partial_{{\nu}}X^{j}\partial^{\underline{\nu}}X^{j}+\frac{1}{2}\bar{\theta}^{I}\Gamma^{\underline{\mu}}\partial_{\mu}\theta^{I}\bar{\theta}^{J}\Gamma^{\underline{\nu}}\partial_{\nu}\theta^{J}+...\ ,$ (3.9)
$L_{2}=$ $-{i}\Big{[}\bar{\theta}^{1}\Gamma^{*}\Gamma^{\underline{\nu}}\partial_{\nu}\theta^{1}-\bar{\theta}^{2}\Gamma^{*}\Gamma^{\underline{\nu}}\partial_{\nu}\theta^{2}\Big{]}+{i}\epsilon^{\mu\nu}\partial_{\mu}X^{i}\Big{[}\bar{\theta}^{1}\Gamma^{i}\partial_{\nu}\theta^{1}-\bar{\theta}^{2}\Gamma^{i}\partial_{\nu}\theta^{2}\Big{]}+...\ ,$ (3.10)
where $\Gamma^{*}\equiv\Gamma^{0}\Gamma^{1}$. Defining the projectors $\mathcal{P}_{\pm}=\frac{1}{2}(1\pm\Gamma^{*})$ we can fix the $\kappa$-symmetry gauge as [4]
$\mathcal{P}_{-}\theta^{1}=\mathcal{P}_{+}\theta^{2}=0,\qquad{i.e.}\qquad\mathcal{P}_{+}\theta^{1}=\theta^{1},\quad\mathcal{P}_{-}\theta^{2}=\theta^{2}.$ (3.11)
Then the expansion does not contain cubic $\theta\theta X$ vertices and we get (omitting the explicit factors of projectors in (3.11)) [^10] Adding $L_{2}$ to $L_{1}$ gives gives the standard kinetic terms for $\theta^{1}$ and $\theta^{2}$ with extra $2P_{\pm}$ factors. The expression below is found after rescaling $\theta^{I}\rightarrow{1\over\sqrt{2}}\theta^{I}$ and omitting the explicit projector factors. We also use the flat-space indices as in (2.23).
$L=$ $-\frac{1}{2}\partial_{a}X^{i}\partial^{{a}}X^{i}+i\bar{\theta}^{I}\Gamma^{a}\partial_{{a}}\theta^{I}$
$+\frac{1}{4}\partial_{a}X^{i}\partial_{b}X^{i}\partial^{{b}}X^{j}\partial^{{a}}X^{j}-\frac{1}{8}\partial_{a}X^{i}\partial^{{a}}X^{i}\partial_{b}X^{j}\partial^{{b}}X^{j}$
$-\frac{i}{2}\partial_{{a}}X^{i}\partial_{{b}}X^{i}\,\bar{\theta}^{I}\Gamma^{{a}}\partial^{b}\theta^{I}+\frac{i}{4}\partial_{{b}}X^{i}\partial^{{b}}X^{i}\,\bar{\theta}^{I}\Gamma^{{a}}\partial_{{a}}\theta^{I}$
$-\frac{1}{4}\bar{\theta}^{I}\Gamma_{({a}}\partial_{b)}\theta^{I}\bar{\theta}^{J}\Gamma^{{a}}\partial^{{b}}\theta^{J}+\frac{1}{8}\bar{\theta}^{I}\Gamma^{{a}}\partial_{a}\theta^{I}\bar{\theta}^{J}\Gamma^{{b}}\partial_{b}\theta^{J}+...\ .$ (3.12)
Under the projection (3.11) $\theta^{I}$ have in total 16 real anticommuting components. Thus we get the same number of physical degrees of freedom and the same structure of the quartic interaction terms as in the spinning string case in (2.23). Note that the signs of the fermionic bilinear terms in (2.23) in (3.12) are opposite but they can be made the same by changing a representation of the gamma matrices.
3.3 ${T\bar{T}}$ deformation of free scalar multiplet
Given that the $(\partial X)^{4}$ terms in the expansion of the Nambu action in the static gauge can be viewed [27] as the $T\bar{T}$ deformation [28] of the free massless scalar theory one may wonder if a similar interpretation may apply also to the spinning string Lagrangian in (2.23). Starting with the free Lagrangian for $D-2$ scalar multiplets on a flat 2d background
$\mathcal{L}_{0}=-\frac{1}{2}\partial_{a}X^{i}\partial^{a}X^{i}-i\bar{\psi}^{i}\gamma^{a}\partial_{a}\psi_{i}\ ,$ (3.13)
the traceless symmetric part of the corresponding stress tensor is [^11] As usual, the symmetry and tracelessness of the stress tensor corresponding to (3.13) hold upon the use of the fermionic equations of motion. Here we are considering an off-shell deformation by the symmetric traceless part only, but a deformation by the full stress tensor should be related to the one below by a redefinition of the $\psi^{i}$ field.
$T_{ab}=\partial_{a}X^{i}\partial_{b}X^{i}+i\bar{\psi}^{i}\gamma_{(a}\partial_{b)}\psi^{i}-\frac{1}{2}\eta_{ab}\big{(}\partial_{c}X^{i}\partial^{c}X^{i}+i\bar{\psi}^{i}\gamma^{c}\partial_{c}\psi^{i}\big{)}\ .$ (3.14)
Then the leading order of the $T\bar{T}$ deformation (${\partial\mathcal{L}\over\partial\lambda}={1\over 2}\det T_{ab}$) is proportional to
$\mathcal{L}_{1}={1\over 2}\det T_{ab}=\frac{1}{4}(T^{a}_{a}T^{b}_{b}-T^{a}_{b}T^{b}_{a})=$ $\ \ \frac{1}{4}\partial_{a}X^{i}\partial_{b}X^{i}\partial^{a}X^{j}\partial^{b}X^{j}-\frac{1}{8}\partial_{a}X^{i}\partial^{a}X^{i}\partial_{b}X^{j}\partial^{b}X^{j}$ (3.15)
$+{i\over 2}\partial_{a}X^{i}\partial_{b}X^{i}\bar{\psi}^{j}\gamma^{a}\partial^{b}\psi^{j}-\frac{i}{4}\partial_{a}X^{i}\partial^{a}X^{i}\bar{\psi}^{j}\gamma^{b}\partial_{b}\psi^{j}$
$-\frac{1}{4}\bar{\psi}^{i}\gamma_{a}\partial_{b}\psi^{i}\bar{\psi}^{j}\gamma^{(a}\partial^{b)}\psi^{j}+\ \frac{1}{8}\bar{\psi}^{i}\gamma^{a}\partial_{a}\psi^{i}\bar{\psi}^{j}\gamma^{b}\partial_{b}\psi^{j}\ .$
The sum $\mathcal{L}=\mathcal{L}_{0}+\mathcal{L}_{1}$ is thus equivalent to (2.23). The corresponding action is invariant under the modified global supersymmetry relating $X^{i}$ and $\psi^{i}$ as discussed in Appendix B.
4 One-loop scattering amplitude
Our aim is to compute the 4-scalar scattering amplitude corresponding to the Lagrangian (2.23). We will follow closely the discussion of the bosonic and GS cases in [1, 16, 4]. The amplitude can be written as
$\mathcal{M}_{ij,kl}=A\,\delta_{ij}\delta_{kl}+B\,\delta_{ik}\delta_{jl}+C\,\delta_{il}\delta_{jk}\ ,$ (4.1)
where $A,B$ and $C$ are the annihilation, transmission and reflection coefficients that are functions of $s,t,u$ (see Appendix A for conventions). These can be computed in the inverse string tension $T^{-1}={2\pi\alpha^{\prime}}$ expansion, e.g., $A=A^{(0)}+T^{-1}A^{(1)}+...$.
The tree-level and one-loop diagrams are shown in Figure 1.
$(p_{1},i)$$(p_{3},k)$$(p_{4},l)$$(p_{2},j)$
$(p_{1},i)$$(p_{2},j)$$(p_{3},k)$$(p_{4},l)$$s$-channel$(p,a)$$(q,b)$
$(p_{1},i)$$(p_{2},j)$$(p_{3},k)$$(p_{4},l)$$t$-channel$(p,a)$$(q,b)$
$(p_{1},i)$$(p_{2},j)$$(p_{4},k)$$(p_{3},l)$$u$-channel$(p,a)$$(q,b)$
Figure 1: Diagrams that contribute to tree-level and one-loop $2\rightarrow\,$2 bosonic scattering amplitude.
The tree-level amplitude is the same as in the Nambu action case, i.e. is found from the $(\partial X)^{4}$ terms in (2.23):
$A^{(0)}=-\frac{1}{2}tu,\qquad B^{(0)}=-\frac{1}{2}su,\qquad C^{(0)}=-\frac{1}{2}st.$ (4.2)
The massless 2d kinematics implies $stu=0$ so that choosing $t=0,\ u=-s$ we get $A^{(0)}=C^{(0)}=0$, $B^{(0)}=-\frac{1}{2}s^{2}$.
The bosonic loop contribution to the amplitude found in [1, 16] is given by (assuming dimensional regularization $d=2-2\epsilon$ and setting $\hat{D}\equiv D-2$)
$A_{b}^{(1)}=$ $-\frac{\hat{D}-6}{96\pi}stu\Big{(}\frac{1}{\epsilon}-\gamma+\log 4\pi\Big{)}$
$-\frac{1}{192\pi}\Big{[}(\hat{D}-24)s^{3}+\big{(}\frac{16}{3}\hat{D}+12-2(\hat{D}-6)\log{-s\over\mu^{2}}\big{)}stu+12\big{(}t\log{s\over t}+u\log{s\over u}\big{)}\Big{]}\ ,$ (4.3)
with $B_{b}^{(1)}$ and $C_{b}^{(1)}$ obtained by interchanging $s,t,u$. Specifying to 2d kinematics with $t=0$ gives
$A_{b}^{(1)}=-C_{b}^{(1)}=-\frac{{\hat{D}}-24}{192\pi}s^{3}\ ,\ \ \ \ \ \ \qquad B_{b}^{(1)}=\frac{i}{16}s^{3},$ (4.4)
so that for $D=26$ one has $A^{(1)}_{b}=C^{(1)}_{b}=0$ and $B^{(1)}_{b}=\frac{i}{16}s^{3}$, in agreement with the integrability of underlying critical string theory.
The $\partial X\partial X\psi\partial\psi$ vertex in (2.23) implies the presence of the fermionic loop contribution to the amplitude. The corresponding vertex multiplying $\delta_{ij}\delta_{kl}$ in momentum space may be written as
$V(p_{1},p_{2},p_{3},p_{4})^{\alpha}_{\beta}=$ $\frac{1}{2}\Big{[}\;(p_{1}\cdot p_{2})\;(p_{3}\cdot\gamma)^{\alpha}_{\beta}-(p_{1}\cdot p_{3})\;(p_{2}\cdot\gamma)^{\alpha}_{\beta}-(p_{2}\cdot p_{3})\;(p_{1}\cdot\gamma)^{\alpha}_{\beta}\Big{]}\ ,$ (4.5)
where $\alpha,\beta=1,2$ are 2d spinor indices. For example, the contribution to the $A$-part of the amplitude in the $s$-channel is ($q=p_{1}+p_{2}-p$) [^12] In computing the amplitude with a loop of Majorana fermions we follow the standard prescription (see, e.g., [29]), i.e. do computation for Dirac fermions and multiply the result by ${1\over 2}$. We also drop the contribution of the second $\partial X\partial X\psi\partial\psi$ term in (2.23) that vanishes on the fermionic equation of motion which does not contribute. It can be eliminated by a redefinition of $\psi$, or, equivalently, the corresponding contribution to the fermionic loop reduces to a tadpole diagram that vanishes in dimensional regularization.
[…]
$N_{A,s}=\hat{D}\ {\rm Tr}\Big{[}V(p_{1},p_{2},p,q)\ (p\cdot\gamma)\ V(-p_{3},-p_{4},q,p)\ (q\cdot\gamma)\Big{]}\ ,$ (4.7)
$p=l+x(p_{1}+p_{2}),\qquad q=(1-x)(p_{1}+p_{2})-l,\qquad\Delta=-x(1-x)s-i\varepsilon\ .$ (4.8)
To evaluate the integral, $N_{A,s}$ can be decomposed in terms of powers of $l^{a}$ as
$N_{A,s}=N_{0}+N_{2}l^{2}+N_{ab}l^{a}l^{b}+N_{4}(l^{2})^{2}+N_{abce}l^{a}l^{b}l^{c}l^{e}+M_{ab}l^{2}l^{a}l^{b},$ (4.9)
$N_{0}=-\frac{1}{16}n_{f}\,\hat{D}\,s^{2}(t+u)^{2}(-1+x)^{2}x^{2},\quad N_{2}=-\frac{1}{16}n_{f}\,\hat{D}\,s(t+u)^{2}(-1+x)x,\quad N_{4}=0\ ,$
$N_{ab}=\frac{1}{16}n_{f}\,\hat{D}\,s(1-x)(mp_{1a}p_{3b}+mp_{2a}p_{4b}+np_{2a}p_{3b}+np_{1a}p_{4b}),$
$M_{ab}=\frac{1}{8}n_{f}\,\hat{D}\,(up_{2a}p_{3b}+up_{1a}p_{4b}+tp_{2_{a}}p_{4b}+tp_{1a}p_{3b}),\qquad N_{abce}=-2n_{f}\,\hat{D}\,p_{1a}p_{2b}p_{3c}p_{4e},$
$\qquad m=s+2(u-2s)x,\quad n=s+2(t-2s)x\ .$
Here $\hat{D}=D-2$ is the number of vector components of $\psi^{i}$ running in the loop and $n_{f}=2$ is the number of 2d spinor components. In dimensional regularization ($d=2-2\epsilon$) the integral (4.6) can be written as
[…]
$\quad K_{1}=N_{2}+\frac{1}{2}N_{ab}\eta^{ab}-2N_{4}\Delta-\Delta M_{ab}\eta^{ab}-\frac{1}{4}\Delta N_{abce}(\eta^{ab}\eta^{ce}+\eta^{ac}\eta^{be}+\eta^{ae}\eta^{bc}),$ (4.11)
$\quad K_{0}=\frac{1}{\Delta}N_{0}-N_{2}+\Delta N_{4}-\frac{1}{2}\Delta M_{ab}\eta^{ab}-\frac{1}{4}\Delta N_{abce}(\eta^{ab}\eta^{ce}+\eta^{ac}\eta^{be}+\eta^{ae}\eta^{bc})\ ,$ (4.12)
where $\Delta$ was defined in (4.8). As a result, the fermionic loop contribution to the $s$-channel $A$-amplitude can be written as
$A^{(1)}_{f,s}=-\frac{1}{2}\frac{n_{f}\,\hat{D}\,stu}{192\pi}\Big{(}\frac{1}{\epsilon}-\gamma+\log 4\pi\Big{)}-\frac{1}{4}\frac{n_{f}\hat{D}}{192\pi}\Big{(}s^{3}+\frac{4}{3}stu-2stu\log\frac{-s}{\mu^{2}}\Big{)}.$ (4.13)
Similarly, we get $A^{(1)}_{f,t}=A^{(1)}_{f,u}=0$ so that the total fermionic loop contributions are
$A_{f}^{(1)}=-\frac{1}{2}\frac{n_{f}\,\hat{D}\,}{192\pi}stu\Big{(}\frac{1}{\epsilon}-\gamma+\log 4\pi\Big{)}-\frac{1}{4}\frac{n_{f}\hat{D}}{192\pi}\Big{[}s^{3}+\frac{4}{3}stu-2stu\log\frac{-s}{\mu^{2}}\Big{]},$ (4.14)
$B_{f}^{(1)}=-\frac{1}{2}\frac{n_{f}\,\hat{D}\,}{192\pi}stu\Big{(}\frac{1}{\epsilon}-\gamma+\log 4\pi\Big{)}-\frac{1}{4}\frac{n_{f}\hat{D}}{192\pi}\Big{[}t^{3}+\frac{4}{3}stu-2stu\log\frac{-t}{\mu^{2}}\Big{]},$ (4.15)
$C_{f}^{(1)}=-\frac{1}{2}\frac{n_{f}\,\hat{D}\,}{192\pi}stu\Big{(}\frac{1}{\epsilon}-\gamma+\log 4\pi\Big{)}-\frac{1}{4}\frac{n_{f}\hat{D}}{192\pi}\Big{[}u^{3}+\frac{4}{3}stu-2stu\log\frac{-u}{\mu^{2}}\Big{]}.$ (4.16)
For $t=0$ this gives
$A_{f}^{(1)}=-C_{f}^{(1)}=-\frac{1}{4}\frac{n_{f}\,\hat{D}}{192\pi}\ s^{3}\ ,\qquad\qquad B_{f}^{(1)}=0\ .$ (4.17)
Summing together the bosonic (4) and the fermionic (4.14) loop contributions we find for the total one-loop $A$ amplitude
$A^{(1)}=$ $-\frac{(1+{1\over 4}n_{f})\hat{D}-6}{96\pi}stu\Big{(}\frac{1}{\epsilon}-\gamma+\log 4\pi\Big{)}$
$-\frac{1}{192\pi}\Big{\{}\Big{[}(1+\tfrac{1}{4}n_{f})\,\hat{D}-24\Big{]}s^{3}+\Big{(}\big{(}\tfrac{16}{3}+\tfrac{1}{3}n_{f}\big{)}\hat{D}+12-2\big{[}(1+\tfrac{1}{4}n_{f})\hat{D}-6\big{]}\log\frac{-s}{\mu^{2}}\Big{)}stu$
$\qquad\qquad\ \ +12\Big{(}t\log\frac{s}{t}+u\log\frac{s}{u}\Big{)}tu\Big{\}}\ ,$ (4.18)
with $B^{(1)}$ and $C^{(1)}$ given by interchanging $s,t,u$ as in (4.14)–(4.16).
Specifying to the 2d kinematics with $t=0$ we get the sum of (4.4) and (4.17), i.e.
$A^{(1)}=-C^{(1)}=-\frac{1}{192\pi}\big{[}(1+\tfrac{1}{4}n_{f})\hat{D}-24\big{]}s^{3},\ \ \qquad B^{(1)}=\frac{i}{16}s^{3}\ .$ (4.19)
In the critical spinning string case where $\hat{D}=8$ and $n_{f}=2$ we thus finish with
$A^{(1)}=-C^{(1)}=\frac{1}{16\pi}s^{3},\qquad\qquad B^{(1)}=\frac{i}{16}s^{3}\ .$ (4.20)
The above expressions for the amplitudes (4.18) and (4.19) with $n_{f}=2$ are the same as found for the GS superstring case [4].
The imaginary part of the amplitude given by $B^{(1)}$ does not receive a contribution from the fermionic loop (cf. (4.17)), i.e. it is the same as in the bosonic string case (4.4) and thus is consistent with "pure-phase" structure of the S-matrix consistent with integrability [1]. As discussed in [4], one can make $A^{(1)}=-C^{(1)}=0$ by adding a local counterterm.
The expected agreement between the spinning string and GS superstring amplitudes follows from the close similarity of the corresponding 4-vertices in the Lagrangians in (2.23) and (3.12). In the GS case $\theta^{I}$ subject to (3.11) have 16 independent components. In the spinning string case $\psi^{i}$ has $n_{f}\,\hat{D}=2(D-2)$ components which for $D=10$ gives the same factor of 16 in the fermionic loop. [^13] It is interesting to note that the coefficient of the UV divergent part of (4.18) (corresponding to the total derivative curvature scalar $\sqrt{-g}R^{(2)}$ term evaluated on the induced metric) given for $n_{f}=2$ by ${3\over 2}\hat{D}-6={3\over 2}(D-6)$ is the same as found in the partition function of the spinning string on a general curved 2d background [30].
In the case of the heterotic string we find from (3.5) that the coefficient $q_{s}=n_{f}\hat{D}$ in the fermion loop contributions (4.14)–(4.17) is replaced by $q_{h}={1\over 2}n_{f}(\hat{D}+N)$ where ${1\over 2}$ is due to the chiral projection and $N$ is the number of right internal fermions $\varphi^{\rm r}$. For the critical string case of $\hat{D}=8$ and $n_{f}=2$ we thus get $q_{h}=40$ instead of $q_{s}=16$ in the spinning string and the superstring cases. The coefficient $\hat{D}+{1\over 4}q_{s}-24=-12$ in (4.19) then becomes $\hat{D}+{1\over 4}q_{h}-24=-6$.
Acknowledgements
We thank F. Seibold for useful discussions and sharing a Mathematica code for computing one-loop amplitudes in the GS string case. AAT is grateful to R. Metsaev for important discussions and suggestions. ZW would like to thank H. Jiang, K. Mkrtchyan and D. Zhong for many discussions. We also thank F. Seibold and R. Metsaev for comments on the draft. This work was supported by the STFC grant ST/T000791/1.
Appendix A Conventions and useful relations
For the $2\rightarrow 2$ scattering, the two incoming massless scalars are labeled by $i,j$ with momenta $p_{1},p_{2}$ and two outgoing ones with $k,l$ with momenta $p_{3},p_{4}$, so that in 2d case
$s=-(p_{1}+p_{2})^{2}=-2p_{1}\cdot p_{2},\qquad t=-(p_{1}-p_{3})^{2}=2p_{1}\cdot p_{3},\qquad u=-(p_{1}-p_{4})^{2}=2p_{1}\cdot p_{4}\ ,$
$\qquad p_{1}+p_{2}=p_{3}+p_{4}\ ,\qquad p^{2}_{r}=0\ ,\qquad\ \ \ s+t+u=0\ ,\ \ \ \ \ \ \ \ \ stu=0\ .$ (A.1)
The $\gamma^{a}$ matrices are chosen as
$\gamma^{0}=\Big{(}\begin{array}[]{cc}0&-1\\ 1&0\end{array}\Big{)},\qquad\qquad\gamma^{1}=\Big{(}\begin{array}[]{cc}0&1\\ 1&0\end{array}\Big{)},\qquad\qquad\gamma^{*}=\gamma^{0}\gamma^{1}=\Big{(}\begin{array}[]{cc}-1&0\\ 0&1\end{array}\Big{)}.$ (A.2)
The 2d fermions $\psi$ have spinor components $(\psi_{R},\psi_{L})$ , with $\bar{\psi}=(\psi_{L},-\psi_{R})$. We also define
$\xi^{\pm}=\xi^{0}\pm\xi^{1},\qquad\partial_{\pm}=\frac{1}{2}(\partial_{0}\pm\partial_{1}),\qquad\psi^{\pm}=\psi^{0}\pm\psi^{1}\ ,$ (A.3)
$\psi^{a}\psi^{\prime}_{a}=-\frac{1}{2}(\psi^{+}\psi^{\prime-}+\psi^{-}\psi^{\prime+}),\qquad\qquad\partial^{a}X^{i}\partial_{a}X^{i}=-4\partial_{+}X^{i}\partial_{-}X^{i}.$ (A.4)
One finds the following relations
$\bar{\psi}^{a}\gamma^{0}\partial_{b}\psi_{a}=-\psi_{L}^{a}\partial_{b}\psi_{a,L}-\psi_{R}^{a}\partial_{b}\psi_{a,R}=\frac{1}{2}\psi_{L}^{+}\partial_{b}\psi^{-}_{L}+\frac{1}{2}\psi_{L}^{-}\partial_{b}\psi^{+}_{L}+\frac{1}{2}\psi_{R}^{+}\partial_{b}\psi^{-}_{R}+\frac{1}{2}\psi_{R}^{-}\partial_{b}\psi^{+}_{R}$ (A.5)
$\bar{\psi}^{a}\gamma^{1}\partial_{b}\psi_{a}=\psi_{L}^{a}\partial_{b}\psi_{a,L}-\psi_{R}^{a}\partial_{b}\psi_{a,R}=-\frac{1}{2}\psi_{L}^{+}\partial_{b}\psi^{-}_{L}-\frac{1}{2}\psi_{L}^{-}\partial_{b}\psi^{+}_{L}+\frac{1}{2}\psi_{R}^{+}\partial_{b}\psi^{-}_{R}+\frac{1}{2}\psi_{R}^{-}\partial_{b}\psi^{+}_{R}\ ,$ (A.6)
$\bar{\psi}^{a}\gamma^{b}\partial_{b}\psi_{a}=-2(\psi_{L}^{a}\partial_{-}\psi_{a,L}+\psi_{R}^{a}\partial_{+}\psi_{a,R})=\psi_{L}^{+}\partial_{-}\psi^{-}_{L}+\psi_{L}^{-}\partial_{-}\psi^{+}_{L}+\psi_{R}^{+}\partial_{+}\psi^{-}_{R}+\psi_{R}^{-}\partial_{+}\psi^{+}_{R}\ .$ (A.7)
Appendix B Global supersymmetry of expanded action
The free part of the Lagrangian (2.23) or $\mathcal{L}_{0}$ in (3.13) is invariant under the standard global supersymmetry (cf. (2.2)), i.e. $\delta X^{i}\sim i\bar{\varepsilon}\psi^{i}\,,\ \delta\psi^{i}\sim\partial X^{i}\varepsilon$. Including the quartic interaction terms in (2.23) or $\mathcal{L}_{1}$ in (3.15) one can check that it is possible to deform these transformations by extra $\partial X\partial X\psi\,\varepsilon+\psi\partial\psi\psi\,\varepsilon$ terms so that the total action is again invariant.
Starting with $\mathcal{L}_{1}$ (3.15) with generic coefficients ${\rm v}_{1},...,{\rm v}_{4}$ in front of the last four fermionic terms and writing it in terms of light-cone derivatives and spinor components we get
$\mathcal{L}_{0}=$ $2\partial_{+}X^{i}\partial_{-}X^{i}+2i\psi_{L}^{i}\partial_{-}\psi_{L}^{i}+2i\psi_{R}^{i}\partial_{+}\psi_{R}^{i}\ ,$
$\mathcal{L}_{1}=$ $2\partial_{+}X^{i}\partial_{+}X^{i}\partial_{-}X^{j}\partial_{-}X^{j}+4{\rm v}_{1}\partial_{-}X^{i}\partial_{-}X^{i}\,\psi_{L}^{j}\partial_{+}\psi_{L}^{j}$
$+4{\rm v}_{1}\partial_{+}X^{i}\partial_{+}X^{i}\,\psi_{R}^{j}\partial_{-}\psi_{R}^{j}+8{\rm v}_{3}\psi_{L}^{i}\partial_{+}\psi_{L}^{i}\,\psi_{R}^{j}\partial_{-}\psi_{R}^{j}$ (B.1)
$+\left(4{\rm v}_{1}+8{\rm v}_{2}\right)\,\partial_{+}X^{i}\partial_{-}X^{i}\,\big{(}\psi_{L}^{j}\partial_{-}\psi_{L}^{j}+\psi_{R}^{j}\partial_{+}\psi_{R}^{j}\big{)}+\left(2{\rm v}_{3}+4{\rm v}_{4}\right)\big{(}\psi_{L}^{i}\partial_{-}\psi_{L}^{i}+\psi_{R}^{i}\partial_{+}\psi_{R}^{i}\big{)}^{2}\ .$
Considering the deformation $\delta=\delta_{(0)}+\delta_{(1)}+...$ of the leading-order supersymmetry transformations
$\delta_{(0)}X^{i}=i\varepsilon_{L}\psi_{R}^{i}-i\varepsilon_{R}\psi_{L}^{i},\qquad\delta_{(0)}\psi_{R}^{i}=-\partial_{-}X^{i}\varepsilon_{L},\qquad\delta_{(0)}\psi_{L}^{i}=\partial_{+}X^{i}\varepsilon_{R}\ ,\qquad\ \ \delta_{(0)}\mathcal{L}_{0}=0\ ,$ (B.2)
and demanding the invariance $(\delta_{(0)}+\delta_{(1)}+...)(\mathcal{L}_{0}+\mathcal{L}_{1}+...)=0$, i.e. $\delta_{(0)}\mathcal{L}_{1}+\delta_{(1)}\mathcal{L}_{0}=0\ ,$ one can fix both the coefficients ${\rm v}_{r}$ in (B.1) and in the deformed supersymmetry transformations $\delta_{(1)}X^{i},\delta_{(1)}\psi^{i}$. One solution is [^14] We assume trivial boundary conditions on an infinite line, i.e. drop all boundary terms.
$\qquad\qquad{\rm v}_{1}=\frac{i}{2},\qquad{\rm v}_{2}=-\frac{i}{4},\qquad{\rm v}_{3}=-\frac{1}{4},\qquad{\rm v}_{4}=\frac{1}{8},$ (B.3)
$\delta_{(1)}X^{i}$ $=-i\partial_{+}X^{i}\partial_{-}X^{j}\psi_{R}^{j}\varepsilon_{L}+i\partial_{-}X^{i}\partial_{+}X^{j}\psi_{L}^{j}\varepsilon_{R}+\frac{1}{2}\psi_{R}^{j}\partial_{+}\psi_{R}^{j}\psi_{L}^{i}\varepsilon_{R}-\frac{1}{2}\psi_{L}^{j}\partial_{-}\psi_{L}^{j}\psi_{R}^{i}\varepsilon_{L}\ ,$
$\delta_{(1)}\psi_{L}^{i}$ $=-(\partial_{+}X^{j})^{2}\partial_{-}X^{i}\varepsilon_{R}-i\partial_{-}X^{j}\partial_{+}\psi_{L}^{i}\psi_{R}^{j}\varepsilon_{L}+i\partial_{-}X^{i}\partial_{+}\psi_{L}^{j}\psi_{L}^{j}\varepsilon_{R}\ ,$
$\delta_{(1)}\psi_{R}^{i}$ $=(\partial_{-}X^{j})^{2}\partial_{+}X^{i}\varepsilon_{L}+i\partial_{+}X^{j}\partial_{-}\psi_{R}^{i}\psi_{L}^{j}\varepsilon_{R}-i\partial_{+}X^{i}\partial_{-}\psi_{R}^{j}\psi_{R}^{j}\varepsilon_{L}\ .$ (B.4)
The corresponding deformed supersymmetry algebra closes on the equations of motion.
The values of the coefficients ${\rm v}_{r}$ in (B.3) correspond precisely to the ones in (2.23) or (3.15). In general, the values of ${\rm v}_{2}$ and ${\rm v}_{4}$ that multiply the terms in $\mathcal{L}_{1}$ that are proportional to the fermionic equations of motion (i.e. the fourth and sixth terms in (3.15)) can be changed by a local redefinition of $\psi^{i}$. [^15] Explicitly, given $\bar{\psi}\gamma^{a}\partial_{a}\psi+f(\xi)\bar{\psi}\gamma^{a}\partial_{a}\psi$ (where here $f\sim\partial X\partial X+\psi\partial\psi$) we can eliminate the $f$-term by redefining $\psi\to\psi^{\prime}=(1-{1\over 2}f+...)\psi$.